Two-country trade model: China and US The demand and supply of autos in China are as follows: Dc = 2000 - 0.02 P, Sc 1200 + 0.03 P The demand and supply of autos in US are as follows: Dus = 1800 - 0.02 P, Sus=1400 + 0.03 P Calculate: a. In the absence of trade, what are the equilibrium price and quantity of autos in China and US? Assume there are no transportation costs. With trade, what is the equilibrium price in the international market? How many autos are traded at this price? How many autos are produced b. and consumed in each country with trade? What is the welfare effect of free trade on each country?


1
Expert's answer
2021-06-16T04:26:29-0400

a. In the absence of trade the equilibrium price and quantity of autos in China and US are:

Dc = Sc,

2000 - 0.02 P = 1200 + 0.03 P,

0.05P = 800,

Pc = 16,000,

Qc = 2000 - 0.02×16,000 = 1,680 units.

Dus = Sus,

1800 - 0.02P = 1400 + 0.03P,

0.05P = 400,

Pus = 8,000,

Qus = 1,800 - 0.02×8,000 = 1,640 units.

With trade the equilibrium price in the international market is:

Di = Dc + Dus = 2000 - 0.02P + 1800 - 0.02P = 3,800 - 0.04P,

Si = Sc + Sus = 1200 + 0.03P + 1400 + 0.03P = 2,600 + 0.06P,

Di = Si,

3,800 - 0.04P = 2,600 + 0.06P,

0.1P = 1,200,

Pi = 12,000,

Qi = 3,800 - 0.04×12,000 = 3,320 units.

b. More autos are produced in US and less in China, more autos are consumed in China and less in US with trade. The welfare effect of free trade will be positive in China and negative in US.